If you add the 2 digits of a 2-digit number, their sum will be 6. The present number is 10 times greater as its opposite number. What is the present number?
Answer by venugopalramana(3286)

LET THE DIGIT IN UNITS PLACE OF THE 2 DIGIT NUMBER BE X....SUM OF 2 DIGITS = 6
HENCE NUMBER IN TENS PLACE =6-X.......ITS VALUE IS 10(6-X)+X=60-10X+X=60-9X...
WHEN THE NUMBER IS REVERSED IT WILL BE X IN TENS PLACE AND 6-X IN UNITS PLACE...SO ITS VALUE IS ..10X+6-X=9X+6.....WE ARE GIVEN THAT THE FIRST NUMBER IS 10 TIMES THE SECOND NUMBER...SO
10(9X+6)=60-9X
90X+60=60-9X
90X+9X=60-60
99X=0
X=0
SO THE NUMBER IS 60.....